# Distinguishing Ignorance from Error in LLM Hallucinations

### v1

@@ 4 Detecting Different Types of Hallucinations
Figure 3: 3-way classification results into (i) hallucinations caused by lack of knowledge (HK-), (ii) hallucinations caused despite having knowledge (HK+), and (iii) factually correct examples. We show result using a bad-shot setting (Top) or an Alice-bob setting (Bottom).
4.2 Generalization of WACK hallucinations across hallucination settings
Next, we examine whether the Bad-shot and Alice-Bob synthetic settings are suitable for investigating hallucinations-despite-knowledge (HK+). To this end, we assess the generalization of hallucination detection probes based on these settings. In particular, we evaluate how well a probe trained on examples from the bad-shot setting generalizes to examples obtained with the Alice-Bob setting (Section 2.2). This presents a significant challenge due to the inherent differences in the prompt between the bad-shots and the Alice-Bob-prompt (unlike the experiment in Section 4.1). In this experiment, we evaluate the ability to differentiate between HK+ and factually-correct examples (binary classification), as the prompt settings are only aimed to split the knowledge examples into those two categories.
Figure 4 displays our findings. While changing the setting used for training the probe reduces the results by up to 10%, it still performs above the random baseline. These results suggest some degree of generalization between different settings used to induce hallucinations despite knowledge. This conclusion lends support to the use of such synthetic datasets for studying the phenomenon of hallucinations despite knowledge.
@@ 4.2 Generalization of WACK hallucinations across hallucination settings
Figure 4: Distinguishing factually correct from hallucinations despite knowledge (HK+), when training on examples from either a bad-shot setting or an Alice-Bob setting, and testing on the Alice-Bob setting. While the change of setting reduces accuracy, the probe still performs substantially above a random baseline.
5 Comparing Model-specific and Generic datasets
This section first compares WACK datasets crafted based on different models. Then it evaluates hallucination detection using model-specific vs. generic datasets.
(a) Knowledge similarity between models.
(b) Hallucination similarity under Bad-shots-prompting.
(c) Hallucination similarity under Alice-Bob-prompting.
Figure 5: Knowledge and Hallucination differences on TriviaQA (above the diagonal) and Natural Questions (below the diagonal) between the models.
@@ 5.1 Different models have different knowledge and different hallucinations
To demonstrate the heterogeneity in knowledge and hallucinations across models, we measure the Jaccard similarity (also known as intersection over union) of WACK datasets generated for different models. To compare knowledge of models, we calculate the Jaccard similarity of examples deemed as high-knowledge in two models, following our procedure from Section 2.1. To compare cases of hallucinations despite knowledge (HK+), we calculate the Jaccard similarity of HK+ examples in two models out of the set of examples that both models know.
Figure 5 displays these similarities. Jaccard values range from 0 (completely dissimilar) to 1 (perfect overlap). In Figure 5(a), knowledge similarity for Natural Questions (below the diagonal) is approximately 0.6, indicating significant knowledge divergence between models. For TriviaQA (above the diagonal), models exhibit higher knowledge similarity (around 0.8).
Figures 5(b) and 5(c) reveal that hallucinations in shared knowledge cases are mostly similar (0.8–0.95). However, a 0.1–0.2 difference in similarity scores suggests each model still exhibits unique hallucination patterns. The bad-shot setting shows lower hallucination similarity than the Alice-Bob setting, indicating greater diversity in hallucination patterns for this scenario. These findings underscore the importance of model-specific approaches to hallucination detection and mitigation, as both knowledge bases and hallucination patterns vary across models and datasets.
@@
Figure 6: Distinguishing factually correct from HK+ hallucinations using classifiers trained on generic vs. model-specific datasets.
5.2 Detecting a model’s hallucinations works better with a model-specific dataset
@@
A generic dataset is a labeled dataset that does not account for model-specific hallucinations or knowledge. Using a generic dataset is a common practice in the field of hallucination research for both detection and mitigation (e.g., Li et al. [2023], Chen et al. [2024], Zhang et al. [2024], Marks and Tegmark [2023], Hoscilowicz et al. [2024]). Thus, for comparison, we also create the generic dataset. The typical method for constructing a labeled QA-closed-book dataset involves using a triplet $q$, $a_{g}$, $a_{h}$, where $q$ is a question, $a_{g}$ is the gold answer, and $a_{h}$ is the hallucinated answer. A hallucination example is created by concatenating $a_{h}$ after $q$, while a factually-correct example is formed by appending $a_{g}$ after $q$. This labeling approach is based on the correctness of the answers relative to world knowledge. However, many datasets only include the $a_{g}$ answer, necessitating the creation of $a_{h}$. Following Li et al. [2023], we generate $a_{h}$ by prompting an LLM to produce a plausible yet incorrect answer. See Appendix B for dataset creation details.
The dataset resembles our model-specific datasets but differs in the origin of hallucinations. The generic dataset’s hallucinations are not model-specific and conflate causes like lack of knowledge (HK-) and hallucinations despite knowledge (HK+). In contrast, our model-specific datasets capture hallucinations unique to each model.
@@ Results.
As this work focuses on cases of hallucinations despite knowledge (HK+), we aim to show that the generic dataset is not effective in catching HK+ hallucinations. Thus, we compare probes trained in two binary settings: (1) model-specific setting, separating HK+ from factually correct examples; and (2) generic setting, separating hallucinations from factually correct examples. We test on the model-specific test set of HK+ and factually correct examples. To make the generic and specific datasets more comparable, we add the bad-shots at the start of the prompt in both cases.
@@
As Figure 6, shows classifiers trained on generic datasets (dashed lines) demonstrate varying degrees of effectiveness and are always worse than classifiers trained on model-specific datasets (solid lines). Notably, the model-specific classifiers maintain relatively high accuracy, unlike their generic counterparts. This comparison underscores the advantages of tailoring hallucination detection methods to individual models, suggesting that this approach more effectively captures model-specific nuances and leads to more reliable identification of hallucinations across various models. These results are consistent with related work [CH-Wang et al., 2023] that showed that a generic detector achieves less on specific datasets than training directly on specific datasets. Unlike us, they use a specific hallucination dataset that does not separate the two hallucinations.
(a) TriviaQA
(b) Natural Questions
Figure 7: Comparing HK+ detection before generation using classifiers trained on model-specific pre-hallucination, generic post-hallucination, and model-specific post-hallucination examples.
5.3 Preemptive Hallucination Detection Using Model-Specific Datasets
@@
So far our detection results were obtained with hidden states obtained after the model processed the answer and potentially have already generated a hallucination. A key advantage of model-specific datasets is their ability to detect potential hallucinations preemptively, before they are generated, a feature not possible with generic datasets. This section explores this capability using our WACK dataset (as before using HK+ and factually correct examples), where each example contains only the question $q$ without an attached answer. This approach allows us to analyze the model’s propensity for hallucination based solely on the input query, which is unfeasible with generic datasets as their labeling typically relies on the concatenated answer, providing no signal of potential hallucination prior to response generation.
@@
Figure 7 presents the preemptive hallucination detection results on the TriviaQA and Natural Questions dataset. Model-specific preemptive hallucination detection (solid lines) demonstrates promising results, indicating the models’ ability to anticipate potential hallucinations. In contrast, generic post-hallucination detection (dashed lines) shows random (and even lower than random) performance, suggesting this approach is ineffective for identifying HK+ hallucinations before they are generated. In comparison, model-specific hallucination detection after generation (dotted lines) yields varied outcomes: for the TriviaQA dataset, some layers and models achieve detection rates approaching 60%–70%, while for Natural Questions, the detection rates remain low and close to random. We conclude that post-hallucination settings are not effective for preemptive hallucination detection, further highlighting the benefits of model-specific datasets.
6 Related Work
Our research investigates hallucination types (HK- and HK+) and develops a methodology for constructing HK+ hallucinations. It is related to research on hallucinations and jailbreaking.
Hallucination Detection.
Detecting hallucinations can involve treating the model as a black box, posing questions or sampling its outputs [Gekhman et al., 2023, Pacchiardi et al., 2023, Manakul et al., 2023, Li et al., 2024a]. Another line of work attempts to detect hallucinations, factuality, or answerability by examining the model’s hidden representations, often by training a detection classifier [Burns et al., 2022, He et al., 2023, Rateike et al., 2023, Slobodkin et al., 2023, Azaria and Mitchell, 2023, CH-Wang et al., 2023, Yuksekgonul et al., 2023, Chen et al., 2023, Yin et al., 2024, Levinstein and Herrmann, 2024, Marks and Tegmark, 2023, Li et al., 2023]. Most of this prior work used generic datasets. While we also employ detectors, we focus on model-specific datasets. Some prior work did explore model-specific hallucination datasets and showed their importance [Azaria and Mitchell, 2023, Ji et al., 2024, Cao et al., 2023, CH-Wang et al., 2023]. However, these efforts did not differentiate between the causes of hallucinations (HK- and HK+).
Jailbreaking.
Jailbreaking refers to techniques for causing LLMs to generate unexpected or incorrect answers. For instance, Zhang et al. [2023] demonstrated the snowballing effect, where once the model outputs an it, it is more likely to generate an incorrect explanation for that fact. Additionally, research has shown that a model’s answers can change due to persuasion, long conversations, fantasy settings, LLM personas, and out-of-distribution prompts [Zeng et al., 2024, Li et al., 2024b, Xu et al., 2023, Yao et al., 2023, Nardo, 2023, Joshi et al., 2023, Pacchiardi et al., 2023]. These studies highlight that the correctness of a model’s output depends on many characteristics of the prompt, allowing hallucinations to occur even when the model knows the correct answer. While their work focuses on identifying methods that induce hallucinations, which can lead to HK+ hallucinations, our investigation directly explores the HK+ phenomenon and its relationship to HK-. Additionally, we introduce a method to automatically construct the WACK dataset for further analysis.
@@ 7 Discussion and conclusion
In this work, we emphasize the importance of differentiating between hallucinations caused by lack of knowledge (HK-) and those occurring despite knowledge (HK+). To focus on the latter, we introduced WACK, a method for creating model-specific datasets based on each model’s knowledge and hallucinations. We proposed two settings to induce HK+ hallucinations, dubbed bad-shots and Alice-Bob, and showed some generalization between them. This indicates that these synthetic settings are effective for studying hallucinations despite knowledge. Our findings also reveal that each model possesses different knowledge and exhibits unique hallucination patterns, highlighting the importance of model-specific datasets. Finally, we showed that generic datasets are less effective at detecting model-specific hallucinations compared to our tailored WACK datasets.
@@ 8 Limitations
Our work has a few limitations. While we evaluated three popular models, the patterns may differ in other ones. Additionally, we used only two settings to induce hallucinations given a model’s correct knowledge; there may be many other ways to achieve similar aims. Finally, we only examined the two extremes of the knowledge spectrum, leaving the middle unexplored.
9 Acknowledgement
This research has been supported by an AI Alignment grant from Open Philanthropy, the Israel Science Foundation (grant No. 448/20), the Azrieli Foundation Early Career Faculty Fellowship, and by a grant under a Master Sponsored Research Agreement between the Technion and Google. We also thank Google Cloud for providing us with credits for running experiments on the Google Cloud Platform.

### v2

@@ 3.2 HK+ Settings
Figure 2: 3-way classification results on TrivaQA into (i) hallucinations caused by lack of knowledge (HK-), (ii) hallucinations caused despite having knowledge (HK+), and (iii) factually correct examples. All the detection results are well above the random baseline (dashed line) indicating the possibility of distinguishing HK+ from HK-.
These settings differ in their degree of mildness. In the Truthful setting, the model is directly prompted to answer the question correctly (an error-free setting). The persona setting introduces additional context of story elements that may confuse the model. The Alice-Bob setting increases complexity further by including typos and implying that the answer might be incorrect. Lastly, the Snowballing setting represents the least mild scenario, where the model is explicitly prompted with incorrect answers.
3.3 HK+ is a Broad Phenomenon
Equipped with our process for separating examples of low and high knowledge (Section 3.1) and further encouraging hallucinations for high-knowledge examples (Section 3.2), we create model-specific datasets. To align with prior work (Li et al., 2023; Zhang et al., 2024a; Chen et al., 2023), we conduct our study using two common general English CBQA datasets: TriviaQA (Joshi et al., 2017) and NaturalQuestions (Kwiatkowski et al., 2019). We experiment with three models: Mistral-7B-v0.3 (Jiang et al., 2023), Llama-3.1-8B (Dubey et al., 2024), and Gemma-2-9B (Team et al., 2024). For more details regarding the dataset construction and additional qualitative evaluation, see Appendix B.
Table 2 provides the HK+ percentage statistics. Across all prompt settings, the models generate 4%–24% HK+ hallucinations out of all cases of high knowledge. The Snowballing setting leads to the highest rate of HK+ examples. Yet, HK+ hallucinations occur even in very mild settings like Truthful and Persona, highlighting that despite high knowledge (strong query/gold-answer associations), models fail to employ their knowledge in various contexts.
4 Mitigating and Detecting Different Types of Hallucinations
We first investigate whether mitigation works differently for HK+ and HK- hallucinations. Then, we examine whether the different types of hallucinations are represented differently inside models.
(a) Knowledge similarity between models.
(b) HK+ similarity under Snowballing.
(c) HK+ similarity under Persona.
(d) Mistral HK+ similarity between settings.
Figure 3: High-Knowledge and HK+ differences on TriviaQA (above the diagonal) and Natural Questions (below the diagonal) between the models and settings. Different models have different knowledge and different HK+ examples in shared knowledge cases.
4.1 Prompt Mitigation Only Works for HK+
To demonstrate the necessity of distinguishing between HK- and HK+ hallucinations, we show that each type requires different mitigation techniques. In particular, HK-, caused by lack of knowledge, cannot be mitigated effectively through simple prompting alone; instead, they require an external knowledge source. Importantly, our goal is not to comprehensively evaluate mitigation techniques but rather to provide evidence supporting the need to distinguish between HK+ and HK- hallucinations, as each type likely requires tailored mitigation methods.
To this end, we prepend the following text before each prompt to encourage the model to generate correct responses (Si et al., 2022; Feng et al., 2024; Taveekitworachai et al., 2024): “Generate answers that are entirely factual and precise, regardless of any issues in the text”. (The specific prompt phrasing is not crucial as other prompts work just as well; see Appendix D.)
Results.
Table 3 shows the QA accuracy when generating 5 tokens for 500 random examples of both HK- and HK+ for all settings. [^4] Importantly, we report accuracy when generating a sequence of tokens, which is more realistic than comparing probabilities of correct and incorrect answer choices. Each cell presents the improvement after using the mitigation.
The values for HK+ are significantly higher than those for HK- across all settings, [^5] For HK-, the values are not zero, which we attribute to a wrong categorization of HK- in a small set of cases underscoring the importance of distinguishing between the two cases for mitigation purposes.
4.2 Simple Classifiers can Moderately Distinguish HK+ from HK- from Factually Correct
We next explore whether a classifier can distinguish between hallucinations arising from a model’s lack of knowledge and those that occur even when the model possesses relevant information. Prior work on hallucination detection did not focus on this important aspect (Section 2). This differentiation is crucial for understanding hallucinations’ underlying mechanisms and developing targeted detection and mitigation strategies. To address a comprehensive scenario, we train a detection classifier to differentiate between three labels, where the model: (1) knows the information and does not hallucinate (’factually correct’), (2) knows the information but hallucinates (HK+), and (3) does not know the information and thus hallucinates (HK-). We detect using the model’s inner states to inspect whether the model represents these hallucinations differently. Note that this is a challenging task, as it not only requires distinguishing hallucinations from factually correct responses, but also determining the type of hallucination. This requires an assessment of the model’s knowledge.
In all detection experiments, we randomly select 1000 examples from each label for analysis in each dataset and split them to 70%/30% for training/testing. [^6] If there are fewer than 1000 hallucinations we use all the examples we have. We use a linear (SVM) classifier for detection, as in prior work (Li et al., 2023; CH-Wang et al., 2023). [^7] The inputs to the classifier are normalized vectors of the model’s inner states at the last answer token. The detection results in the main paper are on hidden states from the residual component (after each Transformer block); see Appendix I for similar results on the MLP and Attention components, and Appendix C for additional implementation details.
@@ 4.2 Simple Classifiers can Moderately Distinguish HK+ from HK- from Factually Correct
Figure 4: Distinguishing factually correct from HK+, when training on examples from either a Snowballing setting or an Alice-Bob setting, and testing on the Alice-Bob setting. While the change of setting reduces accuracy, the classifier still performs substantially above a random baseline.
Results.
Figure 2 shows the detection accuracy results for all models and settings on TriviaQA. (Results on Natural Questions are qualitatively similar; Appendix E.) The accuracy at the highest layer is 60%–70% for snowballing and for the other settings around 50%, across three seeds. As expected, the milder settings have lower accuracy than the Snowballing one, but remain well above the random baseline of 33% (dashed line).These results indicate that the models’ inner states contain some information for differentiating between the three cases, and a linear classifier can moderately work. [^8] We observed similar trends in the binary classification case of distinguishing HK+ from HK-; Appendix E. However, more research is needed to better distinguish between these hallucinations to improve mitigation techniques.
5 Analysis of HK+ Hallucinations
In this section, we analyze HK+ and demonstrate that different models, as well as different settings within the same model, exhibit varying HK+ patterns. Next, we show despite this inconsistency, our framework can still be used to investigate this phenomenon, by demonstrating that the different HK+ settings are generalizable.
@@ 5.1 Different Models Have Different Knowledge and Different Hallucinations
To demonstrate the heterogeneity in knowledge and hallucinations across models, we measure the Jaccard similarity (also known as intersection over union) of WACK datasets generated for different models. Jaccard values range from 00 (completely dissimilar) to $1$ (perfect overlap). First, we compare in Figure 3(a) the Jaccard similarity of examples deemed as high-knowledge in two models, following our procedure from Section 3.1. The knowledge similarity for Natural Questions (below the diagonal) is approximately $0.6$, indicating significant knowledge divergence between models. For TriviaQA (above the diagonal), models exhibit a higher similarity in knowledge (around $0.8$).
Next, we compare the Jaccard similarity of HK+ examples between any two models, out of the set of examples that both models know. Figures 3(b) and 3(c) reveal that HK+ examples are very different between models: 0.10.10.10.10.360.360.360.36 Jaccard scores. This difference suggests each model exhibits unique hallucination patterns. Finally, Figure 3(d) shows that different settings also exhibit unique hallucination patterns within the same model. [^9] See Appendix F for additional results showing similar patterns across configurations. However, as shown in Section 5.2 different settings do demonstrate some generalization.
These findings emphasize the importance of model-specific approaches to hallucination detection and mitigation, as both knowledge bases and hallucination patterns vary across models and prompt settings.
5.2 Generalization of HK+ across Settings
Given that HK+ hallucinations differ between different prompt settings, how reliable is the choice of the setting? In this section we examine the generalization of hallucination detection across settings. If a classifier trained on one setting generalizes well to another one, this would suggest that the classifier captures a general property of HK+. Such findings would indicate that these settings are indeed effective for studying the HK+ phenomenon. Concretely, we evaluate the ability to differentiate between HK+ and factually-correct examples (binary classification), as these settings only aim at splitting questions the model knows how to answer into these two categories.
Figure 4 displays the results when testing on the Alice-Bob setting using classifiers trained on either examples from the same setting (solid lines) or from the Snowballing setting (dashed lines). This presents a significant challenge due to the inherent differences between the two settings. While testing on a setting different than the one used for training lowers the results, they are only lower by an average of 5% at the 15th layer and remain above the random baseline, showing some degree of generalization. (See Appendix J for similar results of generalization between other settings.) These positive results lend support to the use of such synthetic datasets for studying the phenomenon of hallucinations despite knowledge.
(a) Gemma.
(b) Llama.
(c) Mistral.
Figure 5: Distinguishing factually correct from HK+ using classifiers trained on generic vs. model-specific datasets on TriviaQA. Accuracy with classifiers trained on specific datasets is significantly higher than generic dataset. Appendix G has similar results on Natural Questions.
@@ 5.2 Generalization of HK+ across Settings
Figure 6: Comparing preemptive HK+ detection using classifiers trained on model-specific pre-hallucination, generic post-hallucination, and model-specific post-hallucination examples on Mistral. Preemptive hallucination is effective using model-specific datasets, while detectors trained on generic datasets are ineffective.
6 Importance of Model-specific Dataset
After demonstrating that HK+ is a broad phenomenon and highlighting its uniqueness across different models, we now show that it can benefit from model-specific datasets. First, we compare HK+ detection using model-specific vs. generic datasets. Next, we show the possibility of detecting HK+ prior to generation. We argue that these results highlight the importance of model-specific datasets for HK+.
6.1 HK+ Detections Improves with a Model-specific Dataset
@@
A generic dataset is a labeled dataset that does not account for model-specific hallucinations or knowledge. Using a generic dataset is a common practice in the field of hallucination research for both detection and mitigation (e.g., Li et al. (2023); Chen et al. (2024); Zhang et al. (2024a); Marks and Tegmark (2023); Hoscilowicz et al. (2024)). The typical method for constructing a labeled CBQA dataset involves using a triplet $(q,a_{g},a_{h})$, where $q$ is a question, $a_{g}$ is the gold answer, and $a_{h}$ is the hallucinated answer. A hallucination example is created by concatenating $a_{h}$ after $q$, while a factually-correct example is formed by appending $a_{g}$ after $q$. This labeling approach is based on the correctness of the answers relative to world knowledge. However, many datasets only include the $a_{g}$ answer, necessitating the creation of $a_{h}$. Following Li et al. (2023), we generate $a_{h}$ by prompting an LLM (Mistral) to produce a plausible yet incorrect answer. See Appendix B for dataset creation details.
@@ Results.
As this work focuses on HK+ cases (of hallucinations despite knowledge), we aim to show that the generic dataset is not effective in catching them. To this end, we compare classifiers trained in two binary settings: (1) model-specific setting, separating HK+ from factually correct examples; and (2) generic setting, separating hallucinations from factually correct examples. We always test on the model-specific test set of HK+ and factually correct examples. To make the generic and specific datasets more comparable, we add the Snowballing setting at the start of the prompt in both cases.
@@
As Figure 5 shows, classifiers trained on hidden states obtained from generic datasets (dashed lines) demonstrate varying degrees of effectiveness and are always worse than classifiers trained on states from model-specific datasets (solid lines), which maintain relatively high accuracy. This comparison stresses the need to tailor hallucination detection to individual models, as it more effectively captures model-specific nuances and leads to more reliable identification of hallucinations per model. These results are consistent with related work (CH-Wang et al., 2023) that showed that a generic detector performs worse on specific datasets than one trained on specific datasets. However, unlike us, prior work used hallucination datasets that do not distinguish between the two hallucination types, HK- and HK+.
6.2 Model-Specific Datasets Enable Preemptive HK+ Detection
@@
Our previous detection results used hidden states obtained after the model generated an answer, potentially already including a hallucination. A key advantage of model-specific datasets is their ability to detect potential hallucinations preemptively, before they are generated, a feature not possible with generic datasets. This section explores this capability using our WACK dataset (as before, with HK+ and factually correct examples), where each example contains only the question $q$ without an appended answer. As a result, the classifier is trained on the internal states of the examples at the last token of the question, rather than the last token of the answer.
@@
Figure 6 displays preemptive hallucination detection results on the TriviaQA and Natural Questions datasets using the snowballing setting on Mistral. (See Appendix H for similar results on other models.) Model-specific preemptive detection (solid lines) demonstrates strong potential, suggesting that models can effectively anticipate hallucinations before they occur. In contrast, generic post-hallucination detection—evaluated at the last token of the answer (dashed lines)—shows random or even below random performance. This indicates that such an approach is ineffective in identifying HK+ before their generation. Note that the generic approach can only be applied post-answer, as the hallucination signal depends on externally added annotations rather than originating from the model itself. In comparison, model-specific hallucination detection applied at the answer’s last token (dotted lines), produces varied outcomes: for the TriviaQA dataset, some layers achieve detection rates of 60%–70%, while for Natural Questions, the detection rates remain low. We conclude that post-hallucination settings are not effective for preemptive hallucination detection, further highlighting the benefits of model-specific datasets.
@@ 7 Discussion and Conclusion
In this work, we demonstrated the prevalence of HK+ across models, datasets, and settings. To induce HK+, we developed a framework to create a labeled, model-specific dataset that categorizes outputs as HK-, HK+, or factually correct. We examined different configurations and observed some generalization between them, indicating that synthetic settings are effective for studying HK+. Importantly, even mild conditions can induce HK+, showing they manifest without flawed prompts. We highlighted the importance of distinguishing between hallucinations caused by a lack of knowledge (HK-) and those occurring despite knowledge (HK+) by showing that mitigation via prompting can only mitigate HK+. We found that each model exhibits distinct knowledge and hallucination patterns, underscoring the need for model-specific datasets. Additionally, we showed that generic datasets are less effective at detecting model-specific hallucinations, and that model-specific datasets enable preemptive detection.
@@ 8 Limitations
Our work has a few limitations. While we evaluated three popular models, the patterns may differ in other ones. Additionally, we used only four settings to induce hallucinations given a model’s correct knowledge; there may be many other ways to achieve similar aims. Finally, we only examined the two extremes of the knowledge spectrum and used one way of quantifying knowledge. Future work can investigate this middle region.
9 Ethics Statement
Our work provides insights into hallucinations that occur despite the model having the necessary knowledge. While this understanding could be misused to make models less reliable, our goal is to advance the study of the hallucination phenomenon to ultimately improve model reliability.